# A Cost-Benefit Analysis of On-Premise Large Language Model Deployment: Breaking Even with Commercial LLM Services

### v1

@@ VI-E Comprehensive Break-Even Analysis Results
By substituting the values from Table II and Table III into Equation (7), we obtain the results presented in Figure 1. A detailed cost breakdown corresponding to these results is provided in Table V. Our break-even analysis across 54 deployment scenarios shows that the economic viability varies widely. This analysis compares nine open-source models with six commercial API services.
TABLE VI: Break-Even Analysis Summary for All Model-API Combinations (months) with Amortized Performance Differences
@@
Small Model Deployment Economics: Small Model Deployment Economics: Small open-source models are highly cost-effective, often break-even in less than 3 months. For example, the EXAONE-4.0-32B breaks even against Claude-4 Opus in 0.3 months, and against other commercial APIs within a favorable 2-3 months. This allows smaller companies to maintain their competitive AI without the recurring costs.
Medium-Scale Deployment Analysis: Medium-Scale Deployment Analysis: For medium models, break-even periods range from 3.8 to 31.2 months. The GLM-4.5-Air model has the shortest payback against Claude-4 Opus at 3.8 months. Llama-3.3-70B against Gemini 2.5 Pro requires the longest breakeven period at 31.2 months. The \$15,000-\$30,000 hardware cost and \$7.92-\$15.84 monthly operating costs create a significant economic barrier for sustained use.
Large Model Economic Challenges: Large-scale models present the most challenging economic conditions, with break-even horizons stretching anywhere from just over 3 months to nearly 9 years. Among them, the Qwen3-235B illustrates the more favorable end of the spectrum, reaching cost recovery in about 3.5 months when compared against a premium service such as Claude-4 Opus. At the opposite extreme, models like Kimi-K2 illustrate the difficulty of competing with aggressively priced providers: when benchmarked against Gemini 2.5 Pro, the time to break even lengthens dramatically, in this case to over 108 months.
VI-F Commercial API Competitive Analysis
The break-even analysis reveals distinct pricing tiers among commercial providers that fundamentally influence deployment decisions.
Premium Tier - Claude-4 Opus: At the top end, Claude-4 Opus is the most expensive service, averaging \$45 per million tokens. Because of this high cost, local deployment becomes attractive across all model sizes. Break-even can occur in as little as 0.3 months for small models and 3.5–6.9 months for large ones.
Competitive Tier - Claude-4 Sonnet, Grok-4: These providers fall into the middle range, with prices between \$3.13 and \$9.00 per million tokens. Against these baselines, break-even stretches from 1.4 months up to 44.1 months. This exerts moderate economic pressure on medium- and large-scale open-source models, yet it does not eliminate the rationale for local deployment.
Cost-Leadership Tier - Gemini 2.5 Pro, GPT-5: Gemini 2.5 Pro and GPT-5 represents the most cost-competitive commercial option. This aggressive pricing extends break-even periods substantially, ranging from 3.0 months for small models to 63.3 months for large models, challenging the economic viability of on-premise deployment.
VI-G Deployment Decision Framework
We analyzed break-even points by both model size and enterprise type. Small, medium, and large organizations have different needs for computing power, regulatory compliance, and financial resources. These factors shape whether on-premise deployment is practical. Table VI reports the quantitative results. The following sections explain what these results mean for organizations of different sizes.
VI-G1 Small Enterprises (SMEs)
@@ VII Conclusion
This study looks at 54 deployment scenarios and offers organizations a clearer sense of how they might approach LLM application. The findings suggest the economics are not as straightforward as they first seem, and some of the usual assumptions about the practicality of on-premise deployment may not hold up.
  • Small-scale deployments can achieve break-even within as little as 0.3 months relative to premium commercial services, and within at most 3 months under less favorable conditions. This makes local deployment far more accessible than often assumed.
  • Medium-scale deployments demonstrate break-even periods ranging from 2.3 to 34.0 months, creating viable alternatives for organizations with sustained volume requirements
  • Large-scale deployments face break-even periods extending to 69.3 months against cost-competitive services, limiting viability to exceptional organizational circumstances
  • Commercial provider pricing heterogeneity creates order-of-magnitude variations in deployment economics, with premium services enabling rapid cost recovery for local alternatives
Strategic Decision Framework: The study groups deployment scenarios into three categories: those that can pay off quickly (within 0–6 months), those that are longer-term bets (6–24 months), and those that are hard to justify economically (more than 24 months). This approach helps organizations compare different options by weighing both break-even numbers and broader strategic needs.
Market Evolution Context: The costs and benefits of deploying large language models are changing quickly, driven by better models, more efficient hardware, and shifting commercial prices. Standard benchmarks—like those from Artificial Analysis—only show a momentary snapshot, and can become outdated as soon as new optimizations or models appear. This means that choosing how to deploy LLMs isn’t a one-time decision, but more like a long-term investment that has to adapt as the technology and market evolve.
At the same time, it’s becoming increasingly practical for smaller organizations to run their own models on-site, reflecting a broader trend of making LLM technology more accessible. Yet, aggressive pricing from commercial providers continues to push costs down, creating pressure on those who want to run models locally.
Future Research Directions: Future work should empirically validate break-even projections through longitudinal studies of real-world deployments. Total cost of ownership modeling needs to be expanded to incorporate operational realities such as staffing, energy consumption, hardware failures, and maintenance overheads. Hybrid deployment paradigms deserve further exploration, balancing economic efficiency with reliability and strategic flexibility. Finally, new benchmarking frameworks should be developed to track performance and cost metrics over time, enabling fairer and more accurate comparison across the rapidly evolving open-source and commercial ecosystems.
By situating break-even analyses within the broader context of evolving benchmarks, infrastructure costs, and deployment strategies, this survey highlights both the opportunities and the challenges of navigating the shifting landscape of LLM economics.

### v2

@@ VI-E Comprehensive Break-Even Analysis Results
TABLE IV: Break-Even Analysis Summary for All Model-API Combinations (months) with Amortized Performance Differences
@@
By substituting the values from Table I and Table II into Equation (7), we obtain the results summarized in Table IV, with cost details in Table III. Our analysis across 54 deployment scenarios—covering nine open-source models and six commercial APIs—shows substantial variation in economic viability across model sizes.
Small Models. Sub-30B models prove highly cost-effective, often reaching break-even within three months. Their low hardware cost enables smaller organizations to sustain competitive AI capabilities without ongoing subscription fees.
Medium Models. Medium-scale deployments require moderate investment but achieve favorable returns over time. They offer a practical balance between performance and cost and are suitable for enterprises with steady, high-volume workloads.
Large Models. Large-scale setups face the steepest economic barriers. Despite competitive accuracy, their high hardware and power demands extend break-even horizons. This challenge makes them viable only for organizations with sustained, large-scale inference needs.
VI-F Deployment Decision Framework
We analyzed break-even points by both model size and enterprise type. Small, medium, and large organizations have different needs for computing power, regulatory compliance, and financial resources. These factors shape whether on-premise deployment is practical. Table IV reports the quantitative results. The following sections explain what these results mean for organizations of different sizes.
VI-F1 Small Enterprises (SMEs)
@@ VII Conclusion
This study evaluates 54 deployment scenarios to clarify the economic trade-offs of on-premise versus API-based LLM use. Results show that deployment economics are highly context-dependent and challenge common assumptions about local feasibility.
  • Small-scale: Break even within 3 months, making local deployment accessible for smaller organizations.
  • Medium-scale: Offer balanced performance and cost, with recovery typically within 6–24 months.
  • Large-scale: Require longer horizons (often beyond 2 years), feasible only for sustained, high-volume workloads.
  • Pricing variability: Differences among providers create large swings in cost-effectiveness, with premium tiers enabling faster local payback.
Strategic Implications. Deployment choices can be grouped into short-term (0–6 months), mid-term (6–24 months), and long-term ($\geq$24 months) investments, which offers organizations a practical lens to align cost recovery with strategic goals. The rapid evolution of models, hardware, and pricing means deployment is a continuous optimization problem rather than a one-time decision.
Future Work. Further research should empirically validate these break-even estimates and extend TCO models to include staffing and maintenance, and explore hybrid paradigms that combine economic efficiency with reliability. Continued benchmarking will be essential to track performance and cost convergence between open and commercial ecosystems.
Overall, this study situates break-even analysis within the broader landscape of evolving LLM technology and highlights both the growing accessibility of local deployment and the ongoing economic tension with cloud-based services.